Operation Summary for March 14, 2011 - March 16, 2011 |
|
Monday March 14 The day shift began with Operations monitoring the Tevatron (TEV) store 8581 (134.8E30) and the Recycler stash (181.1E10), with the Antiproton Source (Pbar) stacking (11.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE experiment off and in access, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations reported that a Recycler kicker had pre-fired. Operators contacted EE Support; they lowered the kicker’s reservoir voltage. During the evening shift, Operations terminated store 8581 at 5:25 PM, put the TEV into HEP shot setup, and established store 8582 at 6:50 PM with an average initial luminosity (AIL) of 340E30. Tuesday March 15 The midnight shift began with Operations monitoring store 8582 (152.89E30) and the stash (171.3E10), and with Pbar stacking (7.7E10). During the midnight shift, Operations monitored the accelerator complex. During the day shift, operators searched and secured MI-12B. Operations terminated store 8582 at 10:05 AM and put the TEV into shot setup. Operations began sending beam to MiniBooNE at 10:15 AM. Operations established store 8583 at noon with an AIL of 356E30. Switchyard personnel conducted 120 GeV studies. During the evening shift, Operations terminated store 8583 at 10:20 PM and prepared D0 for a collision hall access. Controls personnel accessed enclosure B to replace Meson Septa (MSEP) motors. Wednesday March 16 The midnight shift began with the TEV off for a D0 access, with Operations monitoring the stash (322.2E10), with Pbar stacking (9.5E10), and with MiniBooNE taking beam. During the midnight shift, D0 personnel came out of access at 1 AM. Operations disabled the MiniBooNE beam from 1:20 AM to 2:05 AM. Operations put the TEV into shot setup at 2:10 AM and established store 8587 at 3:17 AM with an AIL of 340E30. Operations reported that a Linac high-energy quadrupole bulk power supply had tripped off frequently during the shift. Operators called in a Linac technician and he replaced the power supply for Q51. Operations reported that stacking stopped when the AP0 target dump water pressure decreased; Pbar personnel came in and added water to the system. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/16/2011 email Fermilab |
| Security, Privacy, Legal | |